OpenStudy (anonymous):

How many solutions does the equation y + 12 = y + 10 + 2 have? Zero One Two Infinitely many

pooja195 (pooja195):

m=m <---infinite m=x <---no solution m=1 <--one sol

pooja195 (pooja195):

solve y + 12 = y + 10 + 2

OpenStudy (anonymous):

y+12 = y+10+2 Y+12 =y+12 every time y is a real number both equations are the same

OpenStudy (anonymous):

@Deeezzzz so the answer is what